Hello can someone tell fringe shift in MM experiment. I dont understand the formula

Grade:12th pass

1 Answers

Vikas TU
7 years ago
fringe shift:
For bright fringe the formulae is:
mk = xmd/D
for dark fringe,
(2m+1)/2 = xmd/D
fror m= 0, 1, 2, 3, …...
where k is the wavelngth of the light
D is the distance b/w the slit and screen.
d is width of slit.